What is the total number of moles of water needed to make 2.5 moles of C6H12O6 ?

CO2+H2O=C6H12O6 + O2
1
Expert's answer
2020-04-02T08:57:56-0400

6CO2 + 6H2O = C6H12O6 + 6O2

n(H2O) = 6 × 2.5 = 15 mol.